The wave function of the hydrogen atom in its ground state (1s) is ψ=(πa03)−12exp(−ra0 where a0=0.529×10−8 cm. The charge density is ρ(x,y,z)=−e|ψ|2, according to the statistical interpretation of the wave function.
Then the value of ⟨r⟩ is given by

a, 5a0^2

b, 2a0^2

c, 4a0^2

d, 3a0^2
1
Expert's answer
2021-10-10T15:56:55-0400

The wave function

"\psi(r)=\frac{1}{\sqrt{\pi}a_0^{3/2}}\exp(-r/a_0)"

The average value of "r^2" is given by

"\langle r^2\rangle=4\pi \int_0^{\infty}r^2\psi^2r^2dr=\frac{4}{a_0^3}\int_0^{\infty}\exp(-2r/a_0)r^4dr"

"=\frac{4}{a_0^3}*\frac{3a_0^5}{4}=3a_0^2"

Answer: d